An asset used in a four-year project falls in the five-year MACRS class (MACRS Table) for tax purposes. The asset has an acquisition cost of $8,200,000 and will be sold for $1,860,000 at the end of the project.
  
If the tax rate is 40 percent, what is the aftertax salvage value of the asset? (Do not round intermediate calculations and round your answer to 2 decimal places, e.g., 32.16.)
  
Aftertax salvage value            $
Answer #1

Total depreciatiom=$8,200,000*(0.2+0.32+0.192+0.1152)=$6783040

Hence book value as on date of sales=(8,200,000-6783040)=$1416960

Hence gain on sales=(1860000-1416960)=$443040

Hence after tax salvage value=Salesproceeds-(Tax rate*Gain on sales)

=1,860,000-(0.4*443040)=$1,682,784.